a water tank can be filled in 20 hours by an inlet pipe and emptied in 25 hours by an outlet pipe. how long will it take to fill the tank if both pipes are left open?
Answer by stanbon(75887)

a water tank can be filled in 20 hours by an inlet pipe and emptied in 25 hours by an outlet pipe. how long will it take to fill the tank if both pipes are left open?
----------------
Inlet pipe DATA:
Time = 20 hrs/job ; Rate = 1/20 job/hr
--------------
Outlet pipe DATA:
Time = 25 hrs/job ; Rate = 1/25 job/hr
---------------
Together DATA:
Time = x hr/job ; Rate = 1/x job/hr
-----------------
EQUATION:
In rate - out rate = together rate
1/20 + 1/25 = 1/x
Multiply thru by 100x to get:
5x + 4x = 100
9x = 100
x = 100/9 = 11 1/9 hrs
================
Cheers,
Stan H.